sin(x)=8/17 and pi/2 less than x less than pi. Find the exact value of tan(2x)
Answer by stanbon(75887)

sin(x)=8/17 and pi/2 less than x less than pi.
----
x is in QII where x is cos(x) is negative.
------
cos(x) = -sqrt[17^2-8^2]/17 = -15/17
--------
tan(x) = sin(x)/cos(x) = (8/17)/(-15/17) = -8/15
-------------------------------------------------
Find the exact value of tan(2x)
tan(2x) = [2tan(x)]/[1-tan^2(x)] = [2(-8/15)]/[1-(64/225)]
= (-16/15)/[(225-64)/225]
======
= (-16/15)/(161/225)
=======
= (-16*15)/161 = -1.49
------------
Cheers,
Stan H.
=============